Find a centre and the principal axes of the following conicoid. Is it ellip-
soids, hyperboloids or something else?
3x2 + 5y2 + 3z2 + 2yz + 2zx + 2xy - 4x - 8z + 5 = 0
1
Expert's answer
2020-06-11T20:31:50-0400

Let,


"f(x,y,x)=3x^2 + 5y^2 + 3z^2 + 2yz + 2zx + 2xy - 4x - 8z + 5 = 0"

Thus, associated quadratic form is

"Q(x,y,z)=3x^2 + 5y^2 + 3z^2 + 2yz + 2zx + 2xy"

Thus, matrix associated with quadratics form is "A=\begin{bmatrix}
 3 & 1&1\\
 1 & 5&1\\
 1&1&3
\end{bmatrix}"

Note that, the eigenvalues of A is

"\begin{vmatrix}
 3-\lambda & 1&1\\
 1 & 5-\lambda&1\\
 1&1&3-\lambda
\end{vmatrix}=0\implies (\lambda-2)(\lambda-3)(\lambda-6)=0"

"(\lambda_1,\lambda_2,\lambda_3)=(2,3,6)" in which all are positive,thus A is positive definite ,thus Hessian matrix "H=2A" is also positive definite.Hence "f(x,y,z)" has one critical point and which is center of the conic and can be found by


"Hx=-p"

where, "p=\begin{bmatrix}
 -4 \\
 0\\-8
\end{bmatrix}" "\implies x=-H^{-1}p" ,On plugin the values we get, center

"x=(\frac{1}{3},-\frac{1}{3},\frac{4}{3})^T"


Now, corresponding to each eigenvalues of A , eigen vectors is

"Av_1=\lambda_1v_1\\
Av_2=\lambda_1v_2\\
Av_3=\lambda_1v_3\\"

On solving we get,

"v_1=(-1,0,1)^T\\
v_2=(1,-1,1)^T\\
v_3=(1,2,1)^T"

Thus ,respective unit vector is,

"\hat{v_1}=\frac{1}{\sqrt{2}}(-1,0,1)^T\\
\hat{v_2}=\frac{1}{\sqrt{3}}(1,-1,1)^T\\
\hat{v_3}=\frac{1}{\sqrt{6}}(1,2,1)^T"

Hence, principle axis, "l_1,l_2,l_3" will be

"l_1=\{(\frac{1}{3},-\frac{1}{3},\frac{4}{3})^T+t\frac{1}{\sqrt{2}}(-1,0,1)^T|t\in \mathbb{R}\}\\
l_2=\{(\frac{1}{3},-\frac{1}{3},\frac{4}{3})^T+t\frac{1}{\sqrt{3}}(1,-1,1)^T|t\in \mathbb{R}\}\\
l_3=\{(\frac{1}{3},-\frac{1}{3},\frac{4}{3})^T+t\frac{1}{\sqrt{6}}(1,2,1)^T|t\in \mathbb{R}\}\\"

Moreover, consider the orthogonal matrix

"\Gamma=(v_1,v_2,v_3)"

And, orthogonal transformation

"x'=\Gamma^Tx"

where,"x=(x_1,y_1,z_1)\&x'=(x'_1,y'_1,z'_z)"

Thus,

"Q(x')=x'^T\Gamma^TA\Gamma x=2x'^2+3y'^2+6z'^2"

Let,

"P(x,y,z)=- 4x - 8z \&r=5"

Thus,

"P(x')=P(\Gamma x)"

Hence, "f(x)" transforms to "Q(x')+P(x')+r=0" .

Now, on simplification we get,

"2X^2+3Y^2+6Z^2=c"

Where,

"X=x'-\frac{1}{\sqrt{2}}\\
Y=y'-\frac{2}{\sqrt{3}}\\
Z=z'-\frac{1}{\sqrt{6}}\\"

and "c>0" a constant.

Therefore, given conicoid is an ellipsoid